# Distributed Coordination for Heterogeneous Non-Terrestrial Networks

### v1

@@ Abstract
To guarantee global coverage and ubiquitous connectivity, the Non-terrestrial Network (NTN) technology has been regarded as a key enabling technology in the Six Generation (6G) network, which consists of the unmanned aerial vehicle (UAV), high-altitude platform (HAP), and satellite. It is noted that the unique characteristics of various NTN platforms directly impact the design and implementation of NTNs, which results in highly dynamic and heterogeneous networks. Even within the same tier, such as the space tier, the NTNs are developed based on different platforms including Low Earth Orbit (LEO), Medium Earth Orbit (MEO), and Geostationary Earth Orbit (GEO). Therefore, distributed coordination among heterogeneous NTNs remains an important challenge. Although distributed learning framework finds a wide range of applications by leveraging rich distributed data and computation resources. The explicit and systematic analysis of the individual layers’ challenges, and corresponding distributed coordination solutions in heterogeneous NTNs has not been proposed yet. In this article, we first summarize the unique characteristics of each NTN platform, and analyze the corresponding impact on the design and implementation of the NTN. We then identify the communication challenges of heterogeneous NTNs in individual layers, where the potential coordinated solutions are identified. We further illustrate the multi-agent deep reinforcement learning (MADRL) algorithms tailored for coordinated solutions in heterogeneous NTNs. Last but not least, we present a case study of the user scheduling optimization problem in heterogeneous UAVs-based cellular networks, where the multi-agent deep deterministic policy gradient (MADDPG) technique is developed to validate the effectiveness of distributed coordination in heterogeneous NTNs.
@@ I Introduction
Due to the improved coverage and resilience against terrestrial infrastructure failures, the Non-terrestrial Network (NTN) has been identified as one of the emerging usage scenarios in International Mobile Telecommunications 2030 (IMT-2030) for the Six Generation (6G) network, which has the potential to provide connections in post-disaster scenarios and remote areas [1]. Currently, the research in the NTN community focuses on three platforms including the unmanned aerial vehicle (UAV), high-altitude platform (HAP), and satellites. It is noted that the unique characteristics of each platform (e.g., power supply and loading capability) play a critical role in the design and implementation of the NTN (e.g., transmission power and number of antennas) [2]. Therefore, the coordination of heterogeneous NTNs remains an important challenge to solve.
@@
Specification Tethered UAV Untethered UAV HAP LEO MEO GEO
Loading Capability 20 Kg 2.7 Kg 140 Kg - - -
Power Supply 30 KWh 0.3 KWh 20 KWh - - -
@@
Endurance 24 hours 30 mins 2 months 5-7 years 10-15 years 15-20 years
@@
Network Type 1 Macro 1 Micro/Pico 7 Multiple Macro 61 Multiple Macro 61 Multiple Macro 61 Multiple Macro
@@
Motivated by this, in this paper, we provide a concrete vision of coordination among heterogeneous NTNs with a focus on the corresponding distributed framework. The main contributions of this article are:
@@
    Based on the unique characteristics of each NTN platform, we then provide a comprehensive analysis of challenges and coordinated solutions in heterogeneous NTNs, where the physical layer, MAC layer, network layer, and application layer are considered in Section III.
@@
    To solve the challenges in Section III in a distributed manner, we further present emerging multi-agent deep reinforcement learning (MADRL) algorithms tailored for coordinated heterogeneous NTNs in Section IV, which are divided into delay-tolerant and delay-sensitive functions, respectively.
@@
    To demonstrate the effectiveness of our proposed distributed learning for heterogeneous NTNs, we present a case study of user scheduling optimization over heterogeneous UAV networks and analyze the results, specifically with the multi-agent deep deterministic policy gradient (MADDPG) algorithm in Section V. Finally, we conclude the article in Section VI.
@@ II Unique Characteristics of Non-terrestrial Network Platforms
Fig. 1 shows the typical terrestrial 5G network consisting of a Radio Unit (RU), a Distributed Unit (DU), a Centralized Unit (CU), and the core network, where O-RAN CatA is adopted for the fronthaul split [6]. To achieve the tradeoff between fronthaul bandwidth and coordination, the beamforming procedure, inverse fast Fourier transform (iFFT), and cyclic prefix processing are integrated within the RU, where the user data symbols and precoding weights are transmitted over the fronthaul. In terms of the Size, Weight and Power (SWaP), it is noted that the typical weights of macro massive MIMO RU, and Baseband Unit (BBU) are over 10Kg and 5Kg, respectively, where BBU consists of DU and CU. The average power consumption of RU and CU can reach 1.175KWh and 0.325KWh under a full workload, which can even reach 3.8KWh with the three-sector setting.
@@ II-A1 Untethered UAV
Taking DJI as an example, the flagship UAV platform Matrice 350 RTK achieves a loading capability of 2.7Kg with 31 minutes of flight time, which cannot carry the macro BS equipment discussed above. Although DJI has customized a delivery UAV platform with the 40Kg loading capability (i.e., DJI Flycart30). The flight time is limited to 8 minutes, which fails to guarantee continuous connections to the users. Apart from that, the battery is originally designed to support the UAV flight with a typical capacity 0.3KWh (e.g., TB65 from DJI). Therefore the commercial untethered UAV can only be utilized to implement a single micro or pico network with RU, DU, and CU integrated as a complete payload.
@@ II-A2 Tethered UAV
To overcome the limitations above, the tethered UAV has been regarded as a promising solution, which has been utilized for emergency communication in post-disaster areas. Based on the power supply provided by the tethered system, the tethered UAV achieves much stronger loading capability and endurance time, which is suitable to implement a single macro network, e.g., DG-M30 tethered UAV can carry 20Kg payloads hovering at 200m for 24 hours. More importantly, the optical fiber provided by the tethered system enables flexible communication payload options, including RU, RU/DU, and RU/DU/CU. However, the tethered system limits the mobility of the UAV, which may lead to coverage degradation.
@@ II-B High Altitude Platform
Operated in the stratosphere e.g., 20km, the HAP is capable of flying for several months continuously, and circles at a radius of a few kilometers or less. Taking the HAP from Stratospheric Platforms Ltd as an example, with a wingspan of 60m, it can support a maximum of 140Kg payload and can provide a 20kW power supply. As indicated in 3GPP, the HAP can support the operation of 7 macro networks with a coverage of 100Km, where the communication payloa can be RU/DU or RU/DU/CU [7]. It is noted that 3GPP supports a maximum cell range up to 100km, which means the HAP can achieve direct connectivity to the regular user terminal without modifying the protocols. More importantly, The relative stationary position of the HAP makes it a promising solution to implement wireless backhaul or midhaul transmission based on point-to-point (P2P) LoS communication. However, It is noted that the environmental temperature of the HAP can be as low as -50 to -70 degrees Celsius. Therefore, an additional heating system should also be considered to guarantee the device’s operation in a suitable temperature range.
@@ III Challenges and Coordinated Solutions in Heterogeneous Non-Terrestrial Networks
Physical Layer Initial Access Cell Selection Cell Shaping
@@ III-A Physical Layer
III-A1 Initial Access
The existing initial access procedures are based on the downlink Reference Signal Received Power (RSRP) of the Synchronization Signal Block (SSB), where the UE connects to the cell with the highest RSRP [9]. In traditional terrestrial networks, most of the cellular cells are homogeneous, where the downlink RSRP can be regarded as an effective indicator of transmission performance for cell selection. However, in the heterogeneous NTNs, cells in each tier have distinct communication capabilities, including the number of antennas, bandwidth, traffic loading, etc. Even for the NTNs in the same tier, the BSs may still have different characteristics, e.g., the tethered UAV and untethered UAV, and the visibility time length of different satellites.
@@ III-A1 Initial Access
III-A2 Channel Estimation
Uplink channel estimation has been regarded as an important step in guiding downlink user data transmission, including scheduling in the MAC layer and precoding in the physical layer, where each UE is required to transmit the uplink pilot signal to the NTNs, i.e., Sounding Reference Signal (SRS). If the UE is associated with a higher tier NTN, higher transmission power is usually needed to compensate for the larger path loss. Therefore, in the vertical heterogeneous NTNs, UEs associated with lower-tier NTNs will experience serious pilot interference from the UEs associated with the higher-tier NTNs. Considering the much larger beam footprint in the higher-tier NTNs and limited beamforming capability at the UE, pilot interference will be more serious than that in traditional terrestrial networks.
To solve this challenge, a coordinated pilot resource allocation scheme should be developed to enhance uplink channel estimation accuracy based on the distribution of terrestrial UEs in each cell. It is noted that the periodicity of pilot transmission can be configured to be one of [2,5,10,20,40,80,160,320] ms, which is generally delay-tolerant in most cases.
III-A3 Precoding
Distributed MIMO (DMIMO) technique has been regarded as a promising solution to reduce inter-cell interference based on coherent joint transmission (CJT), which has been proven to benefit terrestrial cell-edge users significantly [10]. In the CJT scheme, a central server is required to obtain the complete wireless channel (i.e., the wireless channel between one UE and multiple cells), and then perform the precoding algorithm to calculate the precoding matrix, such as weighted minimum mean-square error (WMMSE). It is noted that the precoding procedure is delay-sensitive and executed in every slot, e.g., 0.5 ms with 30KHz subcarrier spacing (SCS).
@@ III-A3 Precoding
III-A4 Power Allocation
Power allocation optimization has been regarded as an effective solution to maximize the channel capacity both in single-cell and multiple-cell scenarios. In the heterogeneous NTNs, cross-link interference (CLI) and remote interference (RI) emerge to be two important challenges due to the LoS-dominated aerial channel, where CLI happens between adjacent cells with different TDD frame structures and RI happens between very distant cells even with synchronized TDD frame structures. It is noted that the upper-tier NTNs usually adopt a much higher downlink transmission power than the lower-tier NTNs, and will cause strong CLI and RI to the lower-tier NTNs. To solve the CLI and RI in heterogeneous NTNs, especially for the downlink (DL) to uplink (UL) direction, a coordinated UL-DL power allocation scheme should be developed to enhance the UL signal strength and mitigate DL interference. Similar to precoding, the power allocation procedure is delay-sensitive and needs to be executed in each time slot.
@@ III-B MAC Layer
III-B1 Scheduling
Different from the DMIMO technique to mitigate inter-cell interference in the physical layer, the scheduling optimization has the potential to enhance the performance of cell-edge users by allocating the same resource block to UEs with near-orthogonal channel characteristics. In the heterogeneous NTNs, the multi-tier NTNs will lead to high coverage overlap, where a highly efficient coordinated scheduling algorithm is needed to satisfy latency requirements in each slot. It is noted that the coordinated scheduling algorithm is also important for the CJT scheme in DMIMO, where the participating cells must transmit the user data on the same resource block. Otherwise, the coordinated cells will sacrifice the power consumption without improving the cell-edge UEs in the primary cell. One possible solution is to pre-schedule the UEs in the primary cell, and share the information with coordinated cells for precoding calculation in the next slot.
@@ III-C Network Layer
III-C1 Routing
Due to the mobility in heterogeneous NTNs, it is difficult to maintain stable point-to-point LoS backhaul transmission between the NTN platforms and the core network on the ground. Therefore, the integrated access and backhaul (IAB) technology has been regarded as a promising solution for more robust backhaul transmission [11]. More importantly, the Backhaul Adaptation Protocol (BAP) layer is introduced in DU for routing, which makes it possible to achieve multi-hop backhaul transmission. However, the dynamic topology and heterogeneous communication capabilities make it challenging to find the optimal routing path. Therefore, the coordinated routing algorithm should be developed by considering both the local cell’s status and the neighboring cells’ status. Another challenge in the backhaul routing via IAB is the channel capacity of the aerial communication. Different from traditional access links, the IAB link transmits the cell-level traffic between two NTNs via an access link, which has a much higher capacity requirement. However, the propagation path is usually dominated by LoS in aerial links, which limits the application of spatial multiplexing to enhance channel capacity. One promising solution is the multi-connectivity technology to enable distributed intelligent backhaul traffic allocation on multiple routing paths. The routing table optimization is classified as a delay-tolerant function, which aims to keep the traffic relatively stable over a period of time and collect enough observations before making a new decision.
@@ III-D Application Layer
III-D1 Trajectory Control
Different from the terrestrial network, heterogeneous NTNs are required to be controlled to optimize the trajectories for collaboratively providing optimal coverage in large-scale areas. Although the satellites have fixed orbits based on their altitude, the satellites still have the capability to change the trajectory with a longer response time at a higher cost, e.g., in emergency communications. Therefore, due to the distinct mobility characteristics of heterogeneous NTNs, the multiple-timescale trajectory optimization algorithm should be developed to enable coordination among NTNs for optimal coverage. More importantly, the received quality of control data over wireless backhaul transmission significantly impacts the trajectory update (e.g., latency of control signal), and then degrades the performance of access links to the terrestrial UEs. Therefore, it remains an important challenge to investigate the closely coupled control signal transmission and user data transmission.
IV Distributed Learning for Coordinated Heterogeneous Non-Terrestrial Networks
It is noted that centralized coordination is challenging for heterogeneous NTNs due to dynamic topology and large delay. Therefore, in this section, we further illustrate the emerging MADRL to facilitate distributed coordination above, where the centralized training and decentralized execution (CTDE) framework is adopted due to higher convergence performance [12]. Based on the delay requirements in Table. II, the NTN-based MADRL solutions tailored for the delay-tolerant distributed coordination and delay-sensitive distributed coordination as follows, respectively:

### v2

@@ Abstract
To ensure global coverage and ubiquitous connectivity, the Non-terrestrial Network (NTN) technology has been regarded as a key enabling technology in the Sixth Generation (6G) network, which consists of the unmanned aerial vehicle (UAV), high-altitude platform (HAP), and satellite. Even within the same tier, such as the space tier, the NTNs are developed based on different platforms, including Low Earth Orbit (LEO), Medium Earth Orbit (MEO), and Geostationary Earth Orbit (GEO). It is noted that the unique characteristics of various NTN platforms directly impact the design and implementation of NTNs, resulting in highly dynamic and heterogeneous networks. Therefore, distributed coordination among heterogeneous NTNs remains an important challenge. Although a distributed learning framework finds a wide range of applications by leveraging rich distributed data and computation resources, the explicit and systematic analysis of the individual layers’ challenges, and corresponding distributed coordination solutions in heterogeneous NTNs has not been proposed yet. In this paper, we first summarize the unique characteristics of each NTN platform and analyze the corresponding impact on its design and implementation. We then identify the communication challenges of heterogeneous NTNs in individual layers, where the potential delay-tolerant or delay-sensitive coordinated solutions are proposed accordingly. We further illustrate the multi-agent deep reinforcement learning (MADRL) algorithms tailored for the proposed coordinated solutions. Last but not least, we present a case study of the joint optimization problem on user scheduling and trajectory design in heterogeneous NTN containing tethered-UAV (T-UAV) and untethered-UAV (U-UAV), where the two-timescale multi-agent deep deterministic policy gradient (MADDPG) technique is developed to validate the effectiveness of distributed coordination.
@@ I Introduction
With the improved coverage and resilience against terrestrial infrastructure failures, the Non-terrestrial Network (NTN) has been identified as one of the emerging usage scenarios in International Mobile Telecommunications 2030 (IMT-2030) for the Six Generation (6G) network, which has the potential to provide connections in post-disaster scenarios and remote areas [1]. Currently, the research in the NTN community focuses on three platforms including the unmanned aerial vehicle (UAV), high-altitude platform (HAP), and satellites. It is noted that the unique characteristics of each platform (e.g., power supply and loading capability) play a critical role in the design and implementation of the NTN (e.g., transmission power and number of antennas) [2]. Therefore, the coordination of heterogeneous NTNs remains an important challenge to solve.
@@
Specification Untethered UAV Tethered UAV HAP LEO MEO GEO
Loading Capability 2.7 kg 20 kg 140 kg - - -
Power Supply 0.3 kWh 30 kWh 20 kWh - - -
@@
Endurance 30 mins 24 hours 2 months 5-7 years 10-15 years 15-20 years
@@
Network Type 1 Micro/Pico 1 Macro 7 Multiple Macro 61 Multiple Macro 61 Multiple Macro 61 Multiple Macro
@@
$\bullet$ RU
[…]
Motivated by this, in this paper, we provide a concrete vision of coordination among heterogeneous NTNs with a focus on the corresponding distributed framework. The main contributions of this paper are as follows:
@@
    We provide a comprehensive analysis of challenges and coordinated solutions in heterogeneous NTNs, where the physical layer, MAC layer, network layer, and application layer are considered in Section III.
@@
    We further present emerging multi-agent deep reinforcement learning (MADRL) algorithms tailored for coordinated heterogeneous NTNs in Section IV, which are divided into delay-tolerant and delay-sensitive functions, respectively.
@@
    We present a case study of joint optimization on user scheduling and trajectory design in a heterogeneous NTN network composed of tethered-UAV (T-UAV) and untethered-UAV (U-UAV), where the multi-agent deep deterministic policy gradient (MADDPG) algorithm is utilized to validate the effectiveness of our proposed distributed coordination method. Finally, we conclude the paper in Section VI.
@@ II Unique Characteristics of Non-terrestrial Network Platforms
Fig. 1 shows the typical terrestrial 5G network consisting of a Radio Unit (RU), a Distributed Unit (DU), a Centralized Unit (CU), and the core network, where O-RAN CatA is adopted for the fronthaul split [6]. To achieve the tradeoff between fronthaul bandwidth and coordination, the beamforming procedure, inverse fast Fourier transform (iFFT), and cyclic prefix processing are integrated within the RU, where the user data symbols and precoding weights are transmitted over the fronthaul. In terms of the Size, Weight, and Power (SWaP), it is noted that the typical weights of macro massive MIMO RU, and Baseband Unit (BBU) are over 10 kg and 5 kg, respectively, where BBU consists of DU and CU. The average power consumption of RU and CU can reach 1.175 kWh and 0.325 kWh under a full workload, which can even reach 3.8 kWh with the three-sector setting.
@@ II-A1 Untethered UAV
Taking DJI as an example, the flagship UAV platform Matrice 350 RTK achieves a loading capability of 2.7 kg with 31 minutes of flight time, which cannot carry the macro BS equipment discussed above. Although DJI has customized a delivery UAV platform with a 40 kg loading capability (i.e., DJI Flycart30). The flight time is limited to 8 minutes, which fails to guarantee continuous connections to the users. Apart from that, the battery is originally designed to support the UAV flight with a typical capacity of 0.3 kWh (e.g., TB65 from DJI). Therefore the commercial untethered UAV can only be utilized to implement a single micro or pico network with RU, DU, and CU integrated as a complete payload.
@@ II-A2 Tethered UAV
To overcome the limitations above, the tethered UAV has been regarded as a promising solution, and it has been utilized for emergency communication in post-disaster areas. Based on the power supply provided by the tethered system, the tethered UAV achieves much stronger loading capability and endurance time, which is suitable to implement a single macro network, e.g., DG-M30 tethered UAV can carry 20 kg payloads hovering at 200 m for 24 hours. More importantly, the optical fiber provided by the tethered system enables flexible communication payload options, including RU, RU/DU, and RU/DU/CU. However, the tethered system limits the mobility of the UAV, which may lead to coverage degradation.
@@ II-B High Altitude Platform
Operated in the stratosphere e.g., 20 km, the HAP is capable of flying for several months continuously, and circles at a radius of a few kilometers or less. Taking the HAP from Stratospheric Platforms Ltd as an example, with a wingspan of 60 m, it can support a maximum of 140 kg payload and provide a 20 kW power supply. As indicated in 3GPP, the HAP can support the operation of 7 macro networks with a coverage of 100 km, where the communication payload can be RU/DU or RU/DU/CU [7]. It is noted that 3GPP supports a maximum cell range of 100 km, which means the HAP can achieve direct connectivity to the regular user terminal without modifying the protocols. More importantly, the relative stationary position of the HAP makes it a promising solution for implementing wireless backhaul or midhaul transmission based on point-to-point (P2P) LoS communication. However, it is noted that the environmental temperature of the HAP can be as low as -50 to -70 degrees Celsius. Therefore, an additional heating system should also be considered to guarantee the device’s operation in a suitable temperature range.
@@ III Challenges and Coordinated Solutions in Heterogeneous Non-Terrestrial Networks
Physical Layer
Initial Access Cell Selection Cell Shaping
@@ III-A Physical Layer
Initial Access: The existing initial access procedures are based on the downlink Reference Signal Received Power (RSRP) of the Synchronization Signal Block (SSB), where the UE connects to the cell with the highest RSRP [9]. In traditional terrestrial networks, most of the cellular cells are homogeneous, where the downlink RSRP can be regarded as an effective indicator of transmission performance for cell selection. However, in the heterogeneous NTNs, cells in each tier have distinct communication capabilities, including the number of antennas, bandwidth, traffic loading, etc. Even for the NTNs in the same tier, the BSs may still have different characteristics, e.g., the tethered UAV and untethered UAV, and the visibility time length of different satellites.
@@
Channel Estimation: Uplink channel estimation has been regarded as an important step in guiding downlink user data transmission, including scheduling in the MAC layer and precoding in the physical layer, where each UE is required to transmit the uplink pilot signal to the NTNs, i.e., Sounding Reference Signal (SRS). If the UE is associated with a higher tier NTN, higher transmission power is usually needed to compensate for the larger path loss. Therefore, in the vertical heterogeneous NTNs, UEs associated with lower-tier NTNs will experience serious pilot interference from the UEs associated with the higher-tier NTNs. Considering the much larger beam footprint in the higher-tier NTNs and limited beamforming capability at the UE, pilot interference will be more serious than that in traditional terrestrial networks.
To solve this challenge, a coordinated pilot resource allocation scheme should be developed to mitigate pilot interference and enhance uplink channel estimation accuracy based on the distribution of terrestrial UEs in each cell. It is noted that the periodicity of pilot transmission can be configured to be one of [2,5,10,20,40,80,160,320] ms, which is generally delay-tolerant in most cases.
Precoding: Distributed MIMO (DMIMO) technique has been regarded as a promising solution to reduce inter-cell interference based on coherent joint transmission (CJT), which has been proven to benefit terrestrial cell-edge users significantly [10]. In the CJT scheme, a central server is required to obtain the information about the complete wireless channel (i.e., the wireless channel between one UE and multiple cells) and then perform the precoding algorithm to calculate the precoding matrix, such as weighted minimum mean-square error (WMMSE). It is noted that the precoding procedure is delay-sensitive and executed in every slot, e.g., 0.5 ms with 30 kHz subcarrier spacing (SCS).
@@
Power Allocation: Power allocation optimization has been regarded as an effective solution to maximize the channel capacity both in single-cell and multiple-cell scenarios. In the heterogeneous NTNs, cross-link interference (CLI) and remote interference (RI) emerge as two important challenges due to the LoS-dominated aerial channel, where CLI happens between adjacent cells with different TDD frame structures, and RI happens between very distant cells even with synchronized TDD frame structures. It is noted that the upper-tier NTNs usually adopt a much higher downlink transmission power than the lower-tier NTNs, which will cause strong CLI and RI to the lower-tier NTNs. To solve the CLI and RI in heterogeneous NTNs, especially for the downlink (DL) to uplink (UL) direction, a coordinated UL-DL power allocation scheme should be developed to enhance the UL signal strength and mitigate DL interference. Similar to precoding, the power allocation procedure is delay-sensitive and needs to be executed in each time slot.
@@ III-B MAC Layer
Scheduling: Different from the DMIMO technique to reduce inter-cell interference in the physical layer, the scheduling optimization has the potential to enhance the performance of cell-edge users by allocating the same resource block to UEs with near-orthogonal channel characteristics. In the heterogeneous NTNs, the multi-tier NTNs will lead to high coverage overlap, where a highly efficient coordinated scheduling algorithm is needed to mitigate the more sophisticated inter-cell interference. It is noted that the coordinated scheduling algorithm is also important for the CJT scheme in DMIMO, where the participating cells must transmit the user data on the same resource block. Otherwise, the coordinated cells will sacrifice the power consumption without improving the cell-edge UEs in the primary cell. One possible solution regarding this issue is to pre-schedule the UEs in the primary cell and share the information with coordinated cells for precoding calculation in the next slot.
@@ III-C Network Layer
Routing: Since diverse mobility patterns of satellites, HAPs, and UAVs lead to the dynamic backhaul topologies, it is difficult to maintain stable point-to-point LoS backhaul transmission between the NTN platforms and the core network on the ground. Therefore, the integrated access and backhaul (IAB) technology has been regarded as a promising solution for more robust backhaul transmission [11], which has been further extended to wireless access backhaul (WAB) at 3GPP Release 19. More importantly, the Backhaul Adaptation Protocol (BAP) layer is introduced in DU for routing, which makes it possible to achieve multi-hop backhaul transmission. However, the dynamic topology and heterogeneous communication capabilities make it challenging to find the optimal routing path. Therefore, the coordinated routing algorithm should be developed by considering both the local cell’s status and the neighboring cells’ status. Another challenge in the backhaul routing via IAB is the channel capacity of the aerial communication. Unlike traditional access links, the IAB link transmits cell-level traffic between two NTNs via an access link, which has a much higher capacity requirement. However, the propagation path is usually dominated by LoS in aerial links, which limits the application of spatial multiplexing to enhance channel capacity. One promising solution is the multi-connectivity technology to enable distributed intelligent backhaul traffic allocation on multiple routing paths. The routing table optimization is classified as a delay-tolerant function, which aims to keep the traffic relatively stable over a period of time and collect enough observations before making a new decision.
@@ III-D Application Layer
Trajectory Control: Different from the terrestrial network, heterogeneous NTNs are required to be controlled to optimize the trajectories for collaboratively providing optimal coverage in large-scale areas. Although the satellites have fixed orbits based on their altitude, the satellites still can change their trajectory with a longer response time at a higher cost, e.g., in emergency communications. Therefore, due to the distinct mobility characteristics of heterogeneous NTNs, the multiple-timescale trajectory optimization algorithm should be developed to enable coordination among NTNs for optimal coverage. More importantly, the received quality of control data over wireless backhaul transmission significantly impacts the trajectory update (e.g., latency of control signal), and then degrades the performance of access links to the terrestrial UEs. Therefore, it remains an important challenge to investigate the closely coupled control signal transmission and user data transmission.
IV MADRL-empowered Coordinated Solutions
The MADRL algorithm has been widely considered as a promising technique to enable efficient distributed coordination in heterogeneous NTNs. Among the typical architectures, the centralized training and execution (CTE) framework suffers from high overhead and propagation delays, while the decentralized training and execution (DTE) framework struggles with non-stationarity and poor coordination. In comparison, the centralized training with decentralized execution (CTDE) framework provides a better balance between scalability and coordination [12], which is adopted as the underlying framework for the MADRL algorithm in our following discussion. Based on the delay requirements in the Table. II, the delay-tolerant and delay-sensitive MADRL-empowered distributed coordination solutions are introduced as follows, respectively: